• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:06:12 on June20,2005

    Entry number 146709

    keyword=Shift Summary

    Day Shift June 20, 2005
    John LeRose
    Elena Kuchina
    Yi Qiang

    Program is Production data taking
    08:30 MCC takes beam for Hall B setup. We go to contolled access to replace A1 PMT.
    10:20 Finished with access. Waiting for beam.
    11:00 Still in contolled access. Roman confirms with visual inspection that everything is OK with the Moller target.
    11:46 MCC says they're ready to give us beam.
    12:02 beam is back!
    13:30 adjusted A1 channel 10 voltage to 2400 (was 2513) before run 4740
    15:30 Running spot++ we notice that the BPM's are giving all zeroes all the time. Problem started during run 4741 continued into 4743
    15:55 it appears that the BPM's are back.
    15:56 BPM's went away again.
    15:59 BPM's Back again


    Run start/stop events
    4737 07:48/08:30 1M
    4738 12:03/12:44 1M
    4739 12:44/13:30 1M
    4740 13:31/14:12 1M
    4741 14:12/14:53 1M BPM's OK at beginning but not at end
    4742 14:53/15:38 1M no BPM info
    4743 15:38/16:03 0.5M BPM's back at end